# ResGS: Residual Densification of 3D Gaussian for Efficient Detail Recovery

### v1

@@ Abstract
Recently, 3D Gaussian Splatting (3D-GS) has prevailed in novel view synthesis, achieving high fidelity and efficiency. However, it often struggles to capture rich details and complete geometry. Our analysis highlights a key limitation of 3D-GS caused by the fixed threshold in densification, which balances geometry coverage against detail recovery as the threshold varies. To address this, we introduce a novel densification method, residual split, which adds a downscaled Gaussian as a residual. Our approach is capable of adaptively retrieving details and complementing missing geometry while enabling progressive refinement. To further support this method, we propose a pipeline named ResGS. Specifically, we integrate a Gaussian image pyramid for progressive supervision and implement a selection scheme that prioritizes the densification of coarse Gaussians over time. Extensive experiments demonstrate that our method achieves SOTA rendering quality. Consistent performance improvements can be achieved by applying our residual split on various 3D-GS variants, underscoring its versatility and potential for broader application in 3D-GS-based applications.
Figure 1: We propose ResGS, a pipeline to boost the rendering quality of 3D-GS [26] while improving efficiency. (a) The 3D-GS densification methods, split and clone, rely on threshold-based selection, in our analysis leads to a trade-off. (b) Our ResGS employs a single densification method, residual split, that adaptively addresses under-reconstruction and over-reconstruction. (c) 3D-GS tends to generate Gaussians with sizes in small variation and low rendering quality. (d) From the same initialization, our method can capture fine details and retrieve sufficient geometry effectively while reducing redundancy.
@@ 1 Introduction
Figure 2: An illustration of the limitation of 3D-GS densification method. (a), (b) and (c) represents 3D-GS with a high threshold, 3D-GS with a low threshold, and 3D-GS with densification method replaced as residual split. We provide visualizations of the final rendered image. A high threshold ensures sufficient structural coverage but causes over-reconstruction, leading to blurry rendering. Conversely, a low threshold mitigates blurry rendering but results in under-reconstruction, compromising coverage and creating a difficult trade-off. Our method, however, adaptively tackles both issues without facing this trade-off.
@@
Novel View Synthesis (NVS) has continued to be a hot topic throughout the years. Traditional methods like point clouds and meshes [4, 15, 9] can achieve high rendering speed but often lack fidelity. On the other hand, neural volumetric approaches [33, 1, 34] can achieve high fidelity for novel views but significantly sacrifice rendering efficiency. Recently, 3D-GS [26] has shown both high fidelity and fast rendering speed. Initialized from point clouds generated by Structure-from-Motion (SfM) [42, 41], 3D-GS represents the scene using a set of anisotropic Gaussian ellipsoids, associated with attributes for geometry (mean, scale, rotation) and appearance (opacity, color). The rendering process projects these 3D Gaussians to the camera plane through a tile-based rasterizer, thus achieving fast rendering speed and high fidelity.
@@
However, 3D-GS still suffers from setbacks, often failing to capture fine details and recover complete geometry, as shown by red and orange rectangles in Fig. 1 and Fig. 2. We investigate the underlying reason for these issues and observe that the densification method in 3DGS faces a trade-off. As illustrated in Fig. 1 (a), during densification in 3DGS [26], the selection between split, designed to address over-reconstruction (where Gaussians cover too much area, resulting in a lack of detail), and clone, aimed at under-reconstruction (missing geometry), is based on a predefined threshold $\tau_{s}$. The Gaussians whose scale $s=max(s_{x},s_{y},s_{z})$ is larger than the threshold undergo split, while those below it use clone. As a result, the reconstruction process is highly sensitive to the chosen threshold. As highlighted by the red rectangles in Fig. 2 (a), a high threshold results in insufficient split, which prevents the model from capturing fine details, leading to blurry artifacts. On the other hand, a lower threshold reduces clone operations, leading to missing geometry, as illustrated in the orange boxes in Fig. 2 (b). This dilemma limits the rendering quality. Furthermore, 3D-GS tends to first split and then clone, resulting in objects being fitted with Gaussians of similar scales. As highlighted by the blue rectangles in Fig. 1, 3D-GS fits textureless areas with excessive number of small Gaussians, which is redundant.
To tackle the aforementioned problems, we propose residual split, a novel densification method avoiding a constant scale threshold to decide clone or split. As illustrated in Fig. 1 (b), the core idea is to generate an asymmetric, downscaled replica as a residual for Gaussians undergoing densification. Thus, our method progressively fits the target shape, with Gaussians adaptively capturing details or covering geometry as needed. Furthermore, we propose a coarse-to-fine training pipeline to supplement the progressiveness of our residual split. Specifically, we structure training in stages, using multi-scale images for gradual supervision. This enables our model to focus on recovering the overall structure in the early stages, gradually refining and retrieving fine details in the later stages. We also develop a progressive Gaussian selection scheme for densification, which gradually encourages coarser Gaussians to undergo further refinement.
@@
  • We propose a novel densification method for 3D Gaussians based on a residual strategy to overcome the limitations of fixed threshold for determining split or clone in 3D-GS;
  • We develop a coarse-to-fine training pipeline, ResGS, for progressive guidance to complement residual split;
  • Our method achieves SOTA rendering quality in NVS on three datasets, including Mip-NeRF360 [1], Tanks&Temples [27], and DeepBlending [20]. Furthermore, applying residual split to multiple 3D-GS variants consistently brings performance improvements, demonstrating that our new densification method is more effective than densification with a fixed scale threshold.
@@ 2.1 Novel View Synthesis
Novel view synthesis (NVS) aims to acquire novel view images for a scene through a set of multi-view images. As a pioneer, NeRF [33] introduced an MLP-based implicit representation. More specifically, it uses an MLP to predict the color and opacity of a point in a scene and uses differentiable volume rendering to render images. Due to its photo-realistic quality, it has been extensively studied recently. Many variants have been designed to fit challenging tasks such as representing 4D scenes [36, 17, 35, 14], few-shot reconstruction [7, 24, 6, 53, 29, 8] and anti-aliasing [1, 2]. However, NeRF-based methods often suffer from long training and rendering time. Therefore, many works [37, 38, 21, 22] have been proposed to extend NeRF into real-time rendering. For example, EfficientNeRF [22] introduced valid sampling and pivotal sampling techniques to accelerate NeRF’s rendering during both training and inference phases.
@@
However, the original 3D-GS still has limitations and myriad works have been proposed to increase the render quality and efficiency of 3D-GS. Mip-Splatting [52] introduced a 2D dilation filter and a 3D smoothing filter to mitigate aliasing and artifacts caused by changes in focal length. Scaffold-GS [31] proposed the concept of neural Gaussians which is to tie Gaussians around an anchor and use an MLP to predict its attributes. Thus, the model is more robust to significant view changes and can improve rendering quality. AbsGS [49] discovered that when accumulating the view space gradients used in densification through pixels, there could be a collision, for gradients at various pixels might have different positivities. They applied an absolute function before summing through pixels to mitigate the issue. Pixel-GS [56] analyzed 3D-GS artifacts from the pixel perspective and introduced a method to select Gaussians for densification based on this analysis. GaussianPro [11] introduced a progressive propagation strategy to propagate Gaussians to under-initialized areas. To reduce redundancy and increase rendering quality, Mini-Splatting [13] proposed a densification and simplification algorithm to reorganize Gaussians’ spatial positions.
@@ 2.2 Progressive Training in NVS
Recently, with the prevalence of 3D-GS, many progressive training pipelines have been proposed. Octree-GS [39] proposed an Octree-based 3D-GS approach, and the Octree depth is progressively deepened. FreGS [54] designed a progressive frequency regulation for 3D-GS which is to calculate the loss of rendered images in the frequency domain and gradually add higher frequency components. Not long ago, MVGS [12] introduced a multi-view regulated training strategy that employs an image pyramid to add detailed information progressively. Although they achieved high performance on rendering quality, they require more massive GPU memory and storage than mainstream methods.
In summary, existing works to optimize 3D-GS primarily concentrate on aspects such as supervision [54, 12], selecting Gaussians for densification [49, 56], and modifying the structure [39, 31] of 3D-GS. None of these approaches address the limitations of the 3D-GS densification method directly or propose an improved alternative.
@@ 3.2 ResGS
We propose ResGS, a coarse-to-fine pipeline to boost the rendering quality of 3D-GS with a novel residual densification operation. Fig. 3 shows the overview of our method. We propose a novel densification method, residual split, (Sec. 3.2.1), to tackle the limitations of the 3D-Gs densification method. Furthermore, we introduce a split-stage supervision schedule to add information to the scene gradually (Sec. 3.2.2). We also design a Gaussian selection scheme that progressively encourages densification for coarse Gaussians to enhance details, as described in Sec. 3.2.3.
@@ 3.2.1 Residual Split
The core idea of our densification method is to generate asymmetrically downscaled Gaussians as residuals to supplement those requiring densification. The newly added Gaussian adaptively enhances the reconstructed model, allowing the scene to be progressively refined. The method is named as residual split. As illustrated in Fig. 3 (a), for a to-be-densified Gaussian $G_{i}$, our densification method consists of two steps. Firstly, we add a down-scaled replicate $G_{j}$. Specifically, its opacity $o_{j}$, rotation $\mathbf{R}_{j}$, spherical harmonics coefficients $SH_{j}$ would be same as $G_{i}$, $\boldsymbol{\mu}_{j}$ is sampled using the original 3D Gaussian as a PDF for sampling, and $\mathbf{S}_{j}$ is defined as $\mathbf{S}_{i}$ divided by a set factor $\lambda_{s}$:
@@
Unlike clone and split which involve hardcoded threshold-based selection, residual split can fill in missing geometry and recover details adaptively without requiring a trade-off, as shown in Fig. 2 (c). For areas where details are challenging to recover, our model initially forms a coarse shape. In subsequent iterations, small-scale residual Gaussians are progressively added to refine and reveal finer details. Conversely, in regions lacking geometry, our model adds residual Gaussians to fill in and complete the missing structure. Moreover, our method’s asymmetrical approach with hierarchical Gaussians of varying scales enables efficient shape fitting. As illustrated by the blue rectangles in Fig. 1, this hierarchy allows us to fill textureless regions with a few large, coarse Gaussians, significantly reducing redundancy. Consequently, our approach delivers higher rendering quality while using fewer Gaussians.
@@ 3.2.2 Image Pyramid as Supervision
Reinforcing progressive training in our densification method, we designed a coarse-to-fine supervision schedule that gradually integrates information into the scene. Specifically, based on the original multi-view training pictures, we build an image pyramid $\{IP_{i}\}_{i=1}^{L}$ consisting of $L$ layers. $IP_{i}$ corresponds to the set of images at different viewpoints in layer $i$. $IP_{L}$ is the set of original pictures:
@@
Correspondingly, our training process is divided into $L$ stages. At stage $i$, we train our model under the supervision of $IP_{i}$, increasing the resolution of the training pictures throughout the stages. Due to the variation of frequency bandwidth in different levels in the image pyramid, high-frequency information is gradually added to the scene. Thus, we can force our model to learn coarse and low-frequency features at early stages and construct an overall structure of the whole scene. At later stages, the focus shifts to capturing high-frequency components, carving details above the coarse expression. Through coarse-to-fine supervision, the model can obtain a division of focus throughout the training process and capture more information from the training images.
@@ 3.2.3 Varying Gradient Threshold
To further emphasize the progressiveness of our pipeline, an incremental approach for selecting Gaussians to densify is designed. Specifically, as iterations progress, the model’s focus gradually shifts toward finer details. To support this, we aim to encourage the densification of coarse Gaussians to capture more detail over time. Due to the asymmetry of our residual split, we can expect the generated Gaussians to be finer than the original ones. This allows us to track the relative fineness level of Gaussians, denoted as $l_{i}$. We assign $l_{i}$ of initial Gaussians to 0. If the level of the Gaussian being densified is $l_{i}$, we assign the newly added Gaussian a level of $l_{i}+1$. Thus, as $l_{i}$ gets higher, Gaussians tend to be smaller and finer.
@@ 4.1 Experiment Settings
We trained three versions of our model: two with AbsGS [49], leveraging its enhanced indication of Gaussians for densification—one standard version and one small version with higher gradient thresholds—and a third version using 3D-GS as the baseline. We applied the periodic opacity reduction from AbsGS across all three versions, as it proved to be a more effective alternative to the opacity reset used in 3D-GS. Additionally, we extended this reduction beyond the densification phase, finding it further minimized redundancy without compromising fidelity. Following all 3D-GS-like methods, we trained our model for 30k iterations. The densification stops at 12000 iterations. For the version with AbsGS, $\tau$ was set to $0.00067$, the small version to $0.0016$; for the version with 3D-GS, $\tau$ was set to $0.00028$. $L$ was set to $3$ and $K$ was set to $3$. Our first stage lasts for 2500 iterations, the second stage for 3500 iterations, and the rest iterations are in the last stage. $\alpha$ was set to $2^{\frac{1}{3}}$, $\lambda_{s}$ was set to $1.6$ and $\beta$ was set to $0.3$. The loss function remains the same as in 3D-GS [26], consisting of a combination of $\mathcal{L}_{1}$ and D-SSIM losses. All experiments were done on an NVIDIA RTX3090 GPU.

### v2

@@ Abstract
Recently, 3D Gaussian Splatting (3D-GS) has prevailed in novel view synthesis, achieving high fidelity and efficiency. However, it often struggles to capture rich details and complete geometry.
Our analysis reveals that the 3D-GS densification operation lacks adaptiveness and faces a dilemma between geometry coverage and detail recovery. To address this, we introduce a novel densification operation, residual split, which adds a downscaled Gaussian as a residual. Our approach is capable of adaptively retrieving details and complementing missing geometry. To further support this method, we propose a pipeline named ResGS. Specifically, we integrate a Gaussian image pyramid for progressive supervision and implement a selection scheme that prioritizes the densification of coarse Gaussians over time. Extensive experiments demonstrate that our method achieves SOTA rendering quality. Consistent performance improvements can be achieved by applying our residual split on various 3D-GS variants, underscoring its versatility and potential for broader application in 3D-GS-based applications.
Figure 1: We propose ResGS, a pipeline to boost the rendering quality of 3D-GS [24] while improving efficiency. (a) The 3D-GS densification operations, split and clone, rely on binary selection and cannot adaptively address the challenges they encounter. (b) Our ResGS employs a single densification operation, residual split, that adaptively addresses under-reconstruction and over-reconstruction. (c) 3D-GS tends to generate Gaussians with sizes in small variation and low rendering quality. (d) From the same initialization, our method can capture fine details and retrieve sufficient geometry effectively while reducing redundancy.
@@ 1 Introduction
Figure 2: An illustration of the limitation of split and clone. (a) selects between split and clone based on a high threshold, (b) a low threshold, and in (c) the densification operation is replaced as residual split. We provide visualizations of the final rendered image. A high threshold ensures sufficient structural coverage but causes over-reconstruction, leading to blurry rendering. Conversely, a low threshold mitigates blurry rendering but results in under-reconstruction, compromising coverage and creating a difficult trade-off. Our method, however, adaptively tackles both issues without facing this trade-off.
@@
Novel View Synthesis (NVS) has continued to be a hot topic throughout the years. Traditional methods like point clouds and meshes [4, 14, 9] can achieve high rendering speed but often lack fidelity. On the other hand, neural volumetric approaches [31, 1, 32] can achieve high fidelity for novel views but significantly sacrifice rendering efficiency. Recently, 3D-GS [24] has shown both high fidelity and fast rendering speed for novel view synthesis. However, 3D-GS still suffers from setbacks, often failing to capture fine details and recover complete geometry, as shown by red and orange rectangles in Fig. 1 and Fig. 2.
While various approaches [48, 12, 50, 43, 13] have been proposed to tackle these challenges and have significantly advanced the rendering quality of 3D-GS-based methods to SOTA levels, they have not modified the densification operation of 3D-GS. In this paper, we take a novel approach by revisiting and refining the densification operation to improve reconstruction quality.
@@
The densification operation used in 3D-GS-based methods consists of two operations: split, designed to address over-reconstruction (where Gaussians cover too much area, resulting in a lack of detail), and clone, aimed at under-reconstruction (missing geometry). However, selecting between the two operations is challenging. 3D-GS-based methods rely on a threshold-based decision mechanism. As illustrated in Fig. 1 (a), Gaussians whose scale $s=max(s_{x},s_{y},s_{z})$ is larger than a predefined threshold $\tau_{s}$ undergo split, while those below it use clone. Due to its simplicity, the selected operation might be suboptimal and bring a dilemma in threshold selection. As highlighted by the red rectangles in Fig. 2 (a), a high threshold results in insufficient split, which prevents the model from capturing fine details, leading to blurry artifacts. On the other hand, a low threshold reduces clone operations, leading to missing geometry, as illustrated in the orange boxes in Fig. 2 (b). This dilemma limits the rendering quality. Furthermore, 3D-GS tends to first split and then clone, resulting in objects being fitted with Gaussians of similar scales. As highlighted by the blue rectangles in Fig. 1, 3D-GS fits textureless areas with an excessive number of small Gaussians, which is redundant.
To tackle the aforementioned problems, we propose residual split, a novel densification operation. As illustrated in Fig. 1 (b), the core idea is that for any under-optimized Gaussian, we spawn a downscaled replica as a residual component while reducing the original Gaussian’s opacity. This simultaneously expands spatial coverage and injects finer details, eliminating the split/clone dilemma.
However, directly applying residual split does not produce significant improvements. This is due to the inefficiency of initialization in 3D-GS. At early iterations, the scene is represented by a few coarse Gaussians that face both over-reconstruction and under-reconstruction. Gaussians added through densification are challenging to optimize, as they must address both issues simultaneously. It’s desired to develop an approach that decouples coverage and detail optimization.
Correspondingly, we propose a coarse-to-fine training pipeline, ResGS, that progressively shifts the focus of training, facilitating optimization. Firstly, we structure the training process into stages, employing multi-scale images for gradual supervision, ensuring that the early stages focus on the overall structure and the later stages focus on details. Secondly, we develop a progressive Gaussian selection scheme for densification, which gradually encourages coarser Gaussians to undergo further refinement. Therefore, finer structures are introduced at later iterations, avoiding premature optimization. These two techniques simplify optimization and thereby improve rendering quality.
@@
  • We propose a novel densification operation for 3D Gaussians based on a residual strategy to overcome the limitations of fixed threshold for determining split or clone in 3D-GS;
  • We develop a coarse-to-fine training pipeline, ResGS, equipped with residual split, which incorporates an image-pyramid-based supervision schedule and a Gaussian selection scheme to enhance rendering quality;
  • Our method achieves SOTA rendering quality in NVS on three datasets, including Mip-NeRF360 [1], Tanks&Temples [25], and DeepBlending [18]. Furthermore, applying residual split to multiple 3D-GS variants consistently brings performance improvements, demonstrating that our new densification operation is more effective than densification with a fixed scale threshold.
@@ 2.1 Novel View Synthesis
Novel view synthesis (NVS) aims to acquire novel view images for a scene through a set of multi-view images. As a pioneer, NeRF [31] introduced an MLP-based implicit representation. More specifically, it uses an MLP to predict the color and opacity of a point in a scene and uses differentiable volume rendering to render images. Due to its photo-realistic quality, it has been extensively studied [7, 22, 6, 47, 27, 8, 1, 2]. However, NeRF-based methods often suffer from long training and rendering time. Therefore, many works [33, 34, 19, 20] have been proposed to extend NeRF into real-time rendering.
@@
However, the original 3D-GS still has limitations, and myriad works have been proposed to increase the render quality and efficiency of 3D-GS. Mip-Splatting [46] introduced a 2D dilation filter and a 3D smoothing filter to mitigate aliasing and artifacts caused by changes in focal length. Scaffold-GS [29] proposed the concept of neural Gaussians which is to tie Gaussians around an anchor and use an MLP to predict its attributes. Thus, the model is more robust to significant view changes and can improve rendering quality. AbsGS [43] discovered that when accumulating the view space gradients used in densification through pixels, there could be a collision, for gradients at various pixels might have different positivities. They applied an absolute function before summing through pixels to mitigate the issue. Pixel-GS [50] analyzed 3D-GS artifacts from the pixel perspective and introduced a method to select Gaussians for densification based on this analysis. GaussianPro [11] introduced a progressive propagation strategy to propagate Gaussians to under-initialized areas. To reduce redundancy and increase rendering quality, Mini-Splatting [13] proposed a densification and simplification algorithm to reorganize Gaussians’ spatial positions.
@@ 2.2 Progressive Training in NVS
Recently, with the prevalence of 3D-GS, many progressive training pipelines have been proposed. Octree-GS [35] proposed an Octree-based 3D-GS approach, and the Octree depth is progressively deepened. FreGS [48] designed a progressive frequency regulation for 3D-GS which is to calculate the loss of rendered images in the frequency domain and gradually add higher frequency components.
@@ 3.2 ResGS
We propose ResGS, a coarse-to-fine pipeline to boost the rendering quality of 3D-GS with a novel residual densification operation. Fig. 3 shows the overview of our method. We propose a novel densification operation, residual split, (Sec. 3.2.1), to tackle the limitations of the 3D-GS densification operation. Furthermore, we introduce a split-stage supervision schedule to add information to the scene gradually (Sec. 3.2.2). We also design a Gaussian selection scheme that progressively encourages densification for coarse Gaussians to enhance details, as described in Sec. 3.2.3.
@@ 3.2.1 Residual Split
The core idea of our densification operation is to generate downscaled Gaussians as residuals to supplement those requiring densification. The newly added Gaussian adaptively enhances the reconstructed model, allowing the scene to be refined as needed. The method is named as residual split. As illustrated in Fig. 3 (a), for a to-be-densified Gaussian $G_{i}$, our densification operation consists of two steps. Firstly, we add a down-scaled replicate $G_{j}$. Specifically, its opacity $o_{j}$, rotation $\mathbf{R}_{j}$, spherical harmonics coefficients $SH_{j}$ would be same as $G_{i}$, $\boldsymbol{\mu}_{j}$ is sampled using the original 3D Gaussian as a PDF for sampling, and $\mathbf{S}_{j}$ is defined as $\mathbf{S}_{i}$ divided by a set factor $\lambda_{s}$:
@@
Figure 4: Intermediate results after applying residual split. We saved the Gaussians at 3,000 and 6,000 iterations after applying residual split. As observed, residual split can adaptively handle over-reconstruction and under-reconstruction by refining blurry regions with finer Gaussians while also effectively reconstructing missing structures. Zoom in for better view.
Unlike clone and split which lack adaptiveness, residual split can fill in missing geometry and recover details adaptively without requiring a trade-off, as shown in Fig. 2 (c). We further illustrate this using intermediate results, as shown in Fig. 4. In areas where details are challenging to recover, down-scaled Gaussians of finer granularity are progressively introduced to refine and reveal intricate details. On the other hand, since densification does not change the scale of the original Gaussian, the total coverage is increased after densification, helping to reconstruct missing structures. Moreover, our method’s nonidentical approach with hierarchical Gaussians of varying scales enables efficient shape fitting. As illustrated by the blue rectangles in Fig. 1, this hierarchy allows us to fill textureless regions with a few large Gaussians, significantly reducing redundancy. Consequently, our approach delivers higher rendering quality while using fewer Gaussians.
@@ 3.2.2 Image Pyramid as Supervision
Building on our motivation in Sec. 1, we designed a coarse-to-fine supervision schedule that gradually integrates information into the scene. Specifically, based on the original multi-view training pictures, we build an image pyramid $\{IP_{i}\}_{i=1}^{L}$ consisting of $L$ layers. $IP_{i}$ corresponds to the set of images at different viewpoints in layer $i$. $IP_{L}$ is the set of original pictures:
@@
Correspondingly, our training process is divided into $L$ stages. At stage $i$, we train our model under the supervision of $IP_{i}$, increasing the resolution of the training pictures throughout the stages. Due to the variation of frequency bandwidth in different levels in the image pyramid, high-frequency information is gradually added to the scene. Thus, we can force our model to learn coarse and low-frequency features at early stages and construct an overall structure of the whole scene. At later stages, the focus shifts to capturing high-frequency components, carving details above the coarse expression.
@@ 3.2.3 Varying Gradient Threshold
To further emphasize progressive training, an incremental approach for selecting Gaussians to densify is designed. Specifically, as iterations progress, the model’s focus gradually shifts toward finer details. To support this, we aim to encourage the densification of coarse Gaussians to capture more detail over time. Due to the nonidentical of our residual split, we can expect the generated Gaussians to be finer than the original ones. This allows us to track the relative fineness level of Gaussians, denoted as $l_{i}$. We assign $l_{i}$ of initial Gaussians to 0. If the level of the Gaussian being densified is $l_{i}$, we assign the newly added Gaussian a level of $l_{i}+1$. Thus, as $l_{i}$ gets higher, Gaussians tend to be smaller and finer. Note that $l_{i}$ is only utilized during training and discarded afterward.
@@ 4.1 Experiment Settings
We trained three versions of our model: two with AbsGS [43], leveraging its enhanced indication of Gaussians for densification—one standard version and one small version with higher gradient thresholds—and a third version using 3D-GS as the baseline. We applied the periodic opacity reduction from AbsGS across all three versions, as it proved to be a more effective alternative to the opacity reset used in 3D-GS. Additionally, we extended this reduction beyond the densification phase, finding it further minimized redundancy without compromising fidelity. Following all 3D-GS-like methods, we trained our model for 30k iterations. The densification stops at 12000 iterations. For the version with AbsGS, $\tau$ was set to $0.00067$, the small version to $0.0016$; for the version with 3D-GS, $\tau$ was set to $0.00028$. $L$ was set to $3$ and $K$ was set to $3$. Our first stage lasts for 2500 iterations, the second stage for 3500 iterations, and the rest iterations are in the last stage. A 500-iteration warm-up phase was introduced between stages with densification disabled. $\alpha$ was set to $2^{\frac{1}{3}}$, $\lambda_{s}$ was set to $1.6$ and $\beta$ was set to $0.3$. The loss function remains the same as in 3D-GS [24], consisting of a combination of $\mathcal{L}_{1}$ and D-SSIM losses. All experiments were done on an NVIDIA RTX3090 GPU.